# Hui Xiong2, Bin Yang1

### v1

@@ Appendix C Full Experimental Results
Dataset Metric CATCH Modern iTrans DualTF ATrans DC TsNet Patch DLin NLin AE Ocsvm IF PCA HBOS
CICIDS A-R 0.795 0.697 0.692 0.603 0.528 0.638 0.732 0.716 0.751 0.691 0.629 0.537 0.787 0.601 0.760
R-A-R 0.737 0.637 0.584 0.434 0.392 0.492 0.684 0.650 0.692 0.609 0.574 0.481 0.739 0.478 0.720
V-ROC 0.743 0.649 0.626 0.485 0.408 0.528 0.693 0.662 0.703 0.631 0.586 0.505 0.750 0.496 0.735
CalIt2 A-R 0.838 0.676 0.791 0.574 0.533 0.527 0.771 0.808 0.752 0.695 0.767 0.804 0.775 0.790 0.798
R-A-R 0.854 0.726 0.817 0.666 0.523 0.505 0.807 0.830 0.794 0.743 0.798 0.840 0.818 0.796 0.847
V-ROC 0.848 0.716 0.809 0.630 0.517 0.503 0.796 0.824 0.781 0.728 0.789 0.828 0.802 0.786 0.831
Credit A-R 0.958 0.957 0.934 0.703 0.552 0.504 0.957 0.957 0.954 0.948 0.909 0.953 0.860 0.871 0.951
R-A-R 0.921 0.919 0.882 0.535 0.379 0.483 0.919 0.918 0.914 0.902 0.843 0.914 0.809 0.784 0.904
V-ROC 0.917 0.908 0.881 0.575 0.398 0.483 0.914 0.911 0.908 0.897 0.840 0.905 0.818 0.805 0.896
GECCO A-R 0.970 0.952 0.795 0.714 0.516 0.555 0.954 0.949 0.947 0.936 0.769 0.804 0.619 0.711 0.557
R-A-R 0.990 0.978 0.884 0.725 0.501 0.609 0.977 0.984 0.987 0.977 0.634 0.757 0.680 0.598 0.510
V-ROC 0.987 0.975 0.871 0.717 0.503 0.593 0.974 0.979 0.982 0.971 0.637 0.756 0.677 0.595 0.503
Genesis A-R 0.974 0.676 0.690 0.937 0.947 0.659 0.913 0.685 0.696 0.755 0.931 0.733 0.549 0.815 0.897
R-A-R 0.981 0.727 0.797 0.975 0.976 0.744 0.919 0.737 0.741 0.791 0.917 0.734 0.693 0.822 0.819
V-ROC 0.978 0.729 0.773 0.971 0.970 0.730 0.913 0.728 0.735 0.787 0.916 0.733 0.660 0.816 0.827
MSL A-R 0.664 0.633 0.611 0.576 0.508 0.507 0.613 0.637 0.624 0.592 0.562 0.524 0.524 0.552 0.574
R-A-R 0.747 0.708 0.686 0.662 0.529 0.596 0.701 0.720 0.703 0.681 0.635 0.594 0.575 0.631 0.643
V-ROC 0.735 0.701 0.678 0.652 0.527 0.587 0.692 0.712 0.695 0.672 0.628 0.590 0.571 0.622 0.635
NYC A-R 0.816 0.466 0.640 0.633 0.671 0.549 0.791 0.709 0.768 0.671 0.504 0.456 0.475 0.666 0.446
R-A-R 0.836 0.598 0.697 0.754 0.753 0.524 0.765 0.722 0.762 0.706 0.636 0.612 0.632 0.744 0.545
V-ROC 0.827 0.579 0.689 0.737 0.722 0.530 0.771 0.719 0.760 0.701 0.619 0.589 0.610 0.730 0.536
PSM A-R 0.652 0.593 0.592 0.600 0.514 0.501 0.592 0.586 0.580 0.585 0.650 0.619 0.542 0.648 0.620
R-A-R 0.640 0.588 0.589 0.507 0.453 0.489 0.593 0.586 0.579 0.585 0.587 0.530 0.543 0.584 0.572
V-ROC 0.639 0.589 0.588 0.507 0.451 0.479 0.593 0.585 0.579 0.585 0.589 0.532 0.542 0.585 0.575
SMD A-R 0.811 0.722 0.745 0.631 0.508 0.502 0.727 0.736 0.728 0.738 0.774 0.602 0.664 0.679 0.626
R-A-R 0.800 0.743 0.762 0.594 0.500 0.505 0.747 0.760 0.754 0.762 0.783 0.579 0.679 0.656 0.597
V-ROC 0.797 0.742 0.761 0.592 0.499 0.505 0.746 0.758 0.751 0.760 0.782 0.578 0.678 0.655 0.597
Contextual4.9 A-R 0.921 0.566 0.916 0.660 0.554 0.535 0.919 0.862 0.709 0.545 0.914 0.706 0.840 0.538 0.470
R-A-R 0.859 0.386 0.851 0.504 0.396 0.405 0.856 0.761 0.543 0.391 0.848 0.517 0.736 0.364 0.310
V-ROC 0.854 0.443 0.845 0.548 0.431 0.422 0.850 0.768 0.586 0.437 0.842 0.563 0.743 0.405 0.362
Contextual7.2 A-R 0.899 0.558 0.893 0.536 0.539 0.514 0.897 0.845 0.691 0.515 0.879 0.716 0.802 0.538 0.458
R-A-R 0.818 0.388 0.812 0.361 0.374 0.436 0.816 0.730 0.522 0.360 0.789 0.530 0.674 0.362 0.297
V-ROC 0.822 0.442 0.814 0.424 0.406 0.443 0.819 0.746 0.567 0.409 0.797 0.585 0.693 0.401 0.351
Global4.8 A-R 0.998 0.885 0.981 0.631 0.560 0.515 0.991 0.993 0.981 0.685 0.996 0.998 0.947 0.770 0.609
R-A-R 0.995 0.804 0.963 0.464 0.388 0.383 0.982 0.986 0.962 0.554 0.991 0.995 0.916 0.654 0.454
V-ROC 0.979 0.806 0.952 0.533 0.430 0.402 0.967 0.967 0.942 0.579 0.972 0.973 0.903 0.687 0.489
Global7.2 A-R 0.997 0.861 0.971 0.559 0.568 0.513 0.988 0.992 0.978 0.665 0.996 0.995 0.928 0.745 0.608
R-A-R 0.993 0.772 0.948 0.395 0.403 0.383 0.977 0.984 0.957 0.534 0.992 0.989 0.886 0.618 0.478
V-ROC 0.979 0.772 0.937 0.442 0.439 0.397 0.962 0.966 0.939 0.560 0.972 0.970 0.878 0.653 0.512
Seasonal4.8 A-R 0.999 0.511 0.945 0.865 0.590 0.661 0.958 0.922 0.819 0.613 0.943 0.840 0.921 0.447 0.522
R-A-R 1.000 0.613 0.929 0.972 0.560 0.642 0.933 0.932 0.847 0.676 0.895 0.849 0.898 0.573 0.629
V-ROC 0.999 0.596 0.929 0.950 0.567 0.634 0.932 0.929 0.839 0.668 0.896 0.845 0.897 0.544 0.609
Seasonal7.7 A-R 0.997 0.513 0.946 0.537 0.579 0.628 0.959 0.921 0.827 0.634 0.954 0.817 0.915 0.426 0.509
R-A-R 0.997 0.640 0.946 0.638 0.558 0.623 0.943 0.931 0.860 0.691 0.907 0.844 0.914 0.583 0.615
V-ROC 0.994 0.615 0.945 0.616 0.559 0.611 0.946 0.929 0.851 0.680 0.915 0.838 0.913 0.546 0.602
Shapelet4.9 A-R 0.982 0.527 0.856 0.623 0.532 0.627 0.871 0.812 0.670 0.552 0.857 0.651 0.767 0.490 0.342
R-A-R 0.996 0.608 0.883 0.742 0.608 0.602 0.880 0.851 0.714 0.642 0.854 0.716 0.800 0.624 0.510
V-ROC 0.992 0.599 0.874 0.723 0.594 0.596 0.875 0.844 0.712 0.627 0.851 0.711 0.794 0.601 0.485
Shapelet7.4 A-R 0.957 0.518 0.872 0.523 0.506 0.567 0.882 0.824 0.697 0.573 0.874 0.660 0.728 0.545 0.332
R-A-R 0.991 0.618 0.877 0.624 0.527 0.546 0.885 0.860 0.748 0.640 0.853 0.745 0.780 0.687 0.512
V-ROC 0.984 0.606 0.876 0.603 0.518 0.537 0.885 0.853 0.741 0.633 0.856 0.727 0.771 0.658 0.486
Mixture5.7 A-R 0.906 0.628 0.859 0.524 0.490 0.577 0.870 0.850 0.692 0.648 0.452 0.440 0.881 0.470 0.443
R-A-R 0.910 0.746 0.909 0.649 0.466 0.519 0.910 0.902 0.759 0.738 0.390 0.386 0.876 0.394 0.465
V-ROC 0.900 0.726 0.899 0.623 0.463 0.522 0.902 0.892 0.745 0.722 0.385 0.381 0.877 0.389 0.458
Mixture5.9 A-R 0.879 0.596 0.834 0.523 0.509 0.562 0.845 0.819 0.650 0.636 0.513 0.503 0.874 0.497 0.492
R-A-R 0.887 0.706 0.866 0.630 0.509 0.552 0.871 0.857 0.723 0.728 0.505 0.476 0.878 0.463 0.548
V-ROC 0.880 0.687 0.860 0.610 0.501 0.542 0.865 0.850 0.708 0.710 0.490 0.462 0.876 0.446 0.531
Trend4.8 A-R 0.944 0.705 0.881 0.561 0.502 0.530 0.877 0.868 0.733 0.719 0.477 0.463 0.943 0.525 0.532
R-A-R 0.956 0.876 0.920 0.620 0.502 0.503 0.924 0.929 0.838 0.833 0.465 0.473 0.906 0.471 0.596
V-ROC 0.953 0.843 0.915 0.617 0.504 0.506 0.917 0.920 0.818 0.811 0.449 0.458 0.910 0.460 0.582
Trend7.8 A-R 0.917 0.821 0.827 0.578 0.542 0.501 0.845 0.858 0.800 0.780 0.509 0.488 0.880 0.509 0.530
R-A-R 0.948 0.882 0.884 0.627 0.537 0.504 0.893 0.903 0.842 0.850 0.531 0.509 0.879 0.483 0.625
V-ROC 0.940 0.874 0.872 0.616 0.530 0.502 0.883 0.895 0.835 0.839 0.509 0.490 0.878 0.471 0.603
$1^{st}$ Count 57 0 0 0 0 0 2 0 0 0 0 2 2 0 0

### v3

@@ Appendix C Full Experimental Results
Dataset Metric CATCH Modern iTrans DualTF ATrans DC TsNet Patch DLin NLin AE TFAD Ocsvm IF PCA HBOS
CICIDS A-R 0.795 0.697 0.692 0.603 0.528 0.638 0.732 0.716 0.751 0.691 0.629 0.504 0.537 0.787 0.601 0.760
R-A-R 0.737 0.637 0.584 0.434 0.392 0.492 0.684 0.650 0.692 0.609 0.574 0.498 0.481 0.739 0.478 0.720
V-ROC 0.743 0.649 0.626 0.485 0.408 0.528 0.693 0.662 0.703 0.631 0.586 0.498 0.505 0.750 0.496 0.735
CalIt2 A-R 0.838 0.676 0.791 0.574 0.533 0.527 0.771 0.808 0.752 0.695 0.767 0.504 0.804 0.775 0.790 0.798
R-A-R 0.854 0.726 0.817 0.666 0.523 0.505 0.807 0.830 0.794 0.743 0.798 0.502 0.840 0.818 0.796 0.847
V-ROC 0.848 0.716 0.809 0.630 0.517 0.503 0.796 0.824 0.781 0.728 0.789 0.501 0.828 0.802 0.786 0.831
Credit A-R 0.958 0.957 0.934 0.703 0.552 0.504 0.957 0.957 0.954 0.948 0.909 0.500 0.953 0.860 0.871 0.951
R-A-R 0.921 0.919 0.882 0.535 0.379 0.483 0.919 0.918 0.914 0.902 0.843 0.498 0.914 0.809 0.784 0.904
V-ROC 0.917 0.908 0.881 0.575 0.398 0.483 0.914 0.911 0.908 0.897 0.840 0.498 0.905 0.818 0.805 0.896
GECCO A-R 0.970 0.952 0.795 0.714 0.516 0.555 0.954 0.949 0.947 0.936 0.769 0.499 0.804 0.619 0.711 0.557
R-A-R 0.990 0.978 0.884 0.725 0.501 0.609 0.977 0.984 0.987 0.977 0.634 0.501 0.757 0.680 0.598 0.510
V-ROC 0.987 0.975 0.871 0.717 0.503 0.593 0.974 0.979 0.982 0.971 0.637 0.500 0.756 0.677 0.595 0.503
Genesis A-R 0.974 0.676 0.690 0.937 0.947 0.659 0.913 0.685 0.696 0.755 0.931 0.497 0.733 0.549 0.815 0.897
R-A-R 0.981 0.727 0.797 0.975 0.976 0.744 0.919 0.737 0.741 0.791 0.917 0.499 0.734 0.693 0.822 0.819
V-ROC 0.978 0.729 0.773 0.971 0.970 0.730 0.913 0.728 0.735 0.787 0.916 0.499 0.733 0.660 0.816 0.827
MSL A-R 0.664 0.633 0.611 0.576 0.508 0.507 0.613 0.637 0.624 0.592 0.562 0.500 0.524 0.524 0.552 0.574
R-A-R 0.747 0.708 0.686 0.662 0.529 0.596 0.701 0.720 0.703 0.681 0.635 0.515 0.594 0.575 0.631 0.643
V-ROC 0.735 0.701 0.678 0.652 0.527 0.587 0.692 0.712 0.695 0.672 0.628 0.514 0.590 0.571 0.622 0.635
NYC A-R 0.816 0.466 0.640 0.633 0.671 0.549 0.791 0.709 0.768 0.671 0.504 0.502 0.456 0.475 0.666 0.446
R-A-R 0.836 0.598 0.697 0.754 0.753 0.524 0.765 0.722 0.762 0.706 0.636 0.509 0.612 0.632 0.744 0.545
V-ROC 0.827 0.579 0.689 0.737 0.722 0.530 0.771 0.719 0.760 0.701 0.619 0.509 0.589 0.610 0.730 0.536
PSM A-R 0.652 0.593 0.592 0.600 0.514 0.501 0.592 0.586 0.580 0.585 0.650 0.500 0.619 0.542 0.648 0.620
R-A-R 0.640 0.588 0.589 0.507 0.453 0.489 0.593 0.586 0.579 0.585 0.587 0.499 0.530 0.543 0.584 0.572
V-ROC 0.639 0.589 0.588 0.507 0.451 0.479 0.593 0.585 0.579 0.585 0.589 0.499 0.532 0.542 0.585 0.575
SMD A-R 0.811 0.722 0.745 0.631 0.508 0.502 0.727 0.736 0.728 0.738 0.774 0.500 0.602 0.664 0.679 0.626
R-A-R 0.800 0.743 0.762 0.594 0.500 0.505 0.747 0.760 0.754 0.762 0.783 0.500 0.579 0.679 0.656 0.597
V-ROC 0.797 0.742 0.761 0.592 0.499 0.505 0.746 0.758 0.751 0.760 0.782 0.500 0.578 0.678 0.655 0.597
ASD A-R 0.824 0.692 0.759 0.579 0.506 0.520 0.805 0.760 0.739 0.690 0.704 0.502 0.588 0.618 0.656 0.603
R-A-R 0.861 0.746 0.812 0.624 0.508 0.515 0.835 0.813 0.793 0.741 0.753 0.506 0.612 0.694 0.682 0.624
V-ROC 0.853 0.739 0.804 0.611 0.507 0.516 0.831 0.806 0.784 0.735 0.745 0.506 0.608 0.685 0.674 0.619
Contextual4.9 A-R 0.921 0.566 0.916 0.660 0.554 0.535 0.919 0.862 0.709 0.545 0.914 0.497 0.706 0.840 0.538 0.470
R-A-R 0.859 0.386 0.851 0.504 0.396 0.405 0.856 0.761 0.543 0.391 0.848 0.498 0.517 0.736 0.364 0.310
V-ROC 0.854 0.443 0.845 0.548 0.431 0.422 0.850 0.768 0.586 0.437 0.842 0.501 0.563 0.743 0.405 0.362
Contextual7.2 A-R 0.899 0.558 0.893 0.536 0.539 0.514 0.897 0.845 0.691 0.515 0.879 0.510 0.716 0.802 0.538 0.458
R-A-R 0.818 0.388 0.812 0.361 0.374 0.436 0.816 0.730 0.522 0.360 0.789 0.466 0.530 0.674 0.362 0.297
V-ROC 0.822 0.442 0.814 0.424 0.406 0.443 0.819 0.746 0.567 0.409 0.797 0.472 0.585 0.693 0.401 0.351
Global4.8 A-R 0.998 0.885 0.981 0.631 0.560 0.515 0.991 0.993 0.981 0.685 0.996 0.501 0.998 0.947 0.770 0.609
R-A-R 0.995 0.804 0.963 0.464 0.388 0.383 0.982 0.986 0.962 0.554 0.991 0.463 0.995 0.916 0.654 0.454
V-ROC 0.979 0.806 0.952 0.533 0.430 0.402 0.967 0.967 0.942 0.579 0.972 0.468 0.973 0.903 0.687 0.489
Global7.2 A-R 0.997 0.861 0.971 0.559 0.568 0.513 0.988 0.992 0.978 0.665 0.996 0.499 0.995 0.928 0.745 0.608
R-A-R 0.993 0.772 0.948 0.395 0.403 0.383 0.977 0.984 0.957 0.534 0.992 0.498 0.989 0.886 0.618 0.478
V-ROC 0.979 0.772 0.937 0.442 0.439 0.397 0.962 0.966 0.939 0.560 0.972 0.502 0.970 0.878 0.653 0.512
Seasonal4.8 A-R 0.999 0.511 0.945 0.865 0.590 0.661 0.958 0.922 0.819 0.613 0.943 0.498 0.840 0.921 0.447 0.522
R-A-R 1.000 0.613 0.929 0.972 0.560 0.642 0.933 0.932 0.847 0.676 0.895 0.528 0.849 0.898 0.573 0.629
V-ROC 0.999 0.596 0.929 0.950 0.567 0.634 0.932 0.929 0.839 0.668 0.896 0.522 0.845 0.897 0.544 0.609
Seasonal7.7 A-R 0.997 0.513 0.946 0.537 0.579 0.628 0.959 0.921 0.827 0.634 0.954 0.506 0.817 0.915 0.426 0.509
R-A-R 0.997 0.640 0.946 0.638 0.558 0.623 0.943 0.931 0.860 0.691 0.907 0.532 0.844 0.914 0.583 0.615
V-ROC 0.994 0.615 0.945 0.616 0.559 0.611 0.946 0.929 0.851 0.680 0.915 0.527 0.838 0.913 0.546 0.602
Shapelet4.9 A-R 0.982 0.527 0.856 0.623 0.532 0.627 0.871 0.812 0.670 0.552 0.857 0.502 0.651 0.767 0.490 0.342
R-A-R 0.996 0.608 0.883 0.742 0.608 0.602 0.880 0.851 0.714 0.642 0.854 0.531 0.716 0.800 0.624 0.510
V-ROC 0.992 0.599 0.874 0.723 0.594 0.596 0.875 0.844 0.712 0.627 0.851 0.525 0.711 0.794 0.601 0.485
Shapelet7.4 A-R 0.957 0.518 0.872 0.523 0.506 0.567 0.882 0.824 0.697 0.573 0.874 0.503 0.660 0.728 0.545 0.332
R-A-R 0.991 0.618 0.877 0.624 0.527 0.546 0.885 0.860 0.748 0.640 0.853 0.530 0.745 0.780 0.687 0.512
V-ROC 0.984 0.606 0.876 0.603 0.518 0.537 0.885 0.853 0.741 0.633 0.856 0.527 0.727 0.771 0.658 0.486
Mixture5.7 A-R 0.906 0.628 0.859 0.524 0.490 0.577 0.870 0.850 0.692 0.648 0.452 0.503 0.440 0.881 0.470 0.443
R-A-R 0.910 0.746 0.909 0.649 0.466 0.519 0.910 0.902 0.759 0.738 0.390 0.511 0.386 0.876 0.394 0.465
V-ROC 0.900 0.726 0.899 0.623 0.463 0.522 0.902 0.892 0.745 0.722 0.385 0.510 0.381 0.877 0.389 0.458
Mixture8.9 A-R 0.879 0.596 0.834 0.523 0.509 0.562 0.845 0.819 0.650 0.636 0.513 0.500 0.503 0.874 0.497 0.492
R-A-R 0.887 0.706 0.866 0.630 0.509 0.552 0.871 0.857 0.723 0.728 0.505 0.517 0.476 0.878 0.463 0.548
V-ROC 0.880 0.687 0.860 0.610 0.501 0.542 0.865 0.850 0.708 0.710 0.490 0.516 0.462 0.876 0.446 0.531
Trend4.8 A-R 0.944 0.705 0.881 0.561 0.502 0.530 0.877 0.868 0.733 0.719 0.477 0.502 0.463 0.943 0.525 0.532
R-A-R 0.956 0.876 0.920 0.620 0.502 0.503 0.924 0.929 0.838 0.833 0.465 0.521 0.473 0.906 0.471 0.596
V-ROC 0.953 0.843 0.915 0.617 0.504 0.506 0.917 0.920 0.818 0.811 0.449 0.519 0.458 0.910 0.460 0.582
Trend7.8 A-R 0.917 0.821 0.827 0.578 0.542 0.501 0.845 0.858 0.800 0.780 0.509 0.501 0.488 0.880 0.509 0.530
R-A-R 0.948 0.882 0.884 0.627 0.537 0.504 0.893 0.903 0.842 0.850 0.531 0.528 0.509 0.879 0.483 0.625
V-ROC 0.940 0.874 0.872 0.616 0.530 0.502 0.883 0.895 0.835 0.839 0.509 0.525 0.490 0.878 0.471 0.603
$1^{st}$ Count 65 0 0 1 0 0 1 0 0 0 0 0 0 2 0 0